OpenStudy (anonymous):

I NEED HELP FAST PLEASE WILL FAN AND MEDAL!!!!!!!! Just simple algebra! Express tan(A) in terms of x, y and/or z. Note: z and y are sides of the right triangle and x is the hypotenuse. The right triangle is pointing to the right

OpenStudy (shiraz14):

is A the angle opposite the side labeled x or y?

OpenStudy (anonymous):

choices are y/x z/x z/y y/z and y is on the opposite side of the triangle from angle A

OpenStudy (anonymous):

y is the only vertical line on the triangle

OpenStudy (shiraz14):

tan(A)=y/z

OpenStudy (anonymous):

Thanks you so much! could you help me with one more?

OpenStudy (shiraz14):

You're welcome. And yes, I can.

OpenStudy (anonymous):

The size of a television is given by the length of the diagonal of the screen. The ratio of the height to the width of wide flat-screen TV is 8:15. Find expressions for the width and height of a wide-screen TV in terms of the length of its diagonal

OpenStudy (anonymous):

answer choices: w=15/17 d; h=8/17 d w=8/17 d; h=15/17 d w=15/13 d; h = 8/23 d w = 8/23 d; h=16/23 d

OpenStudy (shiraz14):

You would use Pythagoras' Theorem for this question. Length of diagonal = \[\sqrt{8^{2}+15^{2}}\] = 17 Let x represent the length of the diagonal. Then we have the following: width = 15x/17; height = 8x/17.

OpenStudy (anonymous):

You have been so much of a help thank you! :)

OpenStudy (anonymous):

you have a new fan :)

OpenStudy (anonymous):

Also im going to put up another question but not in here because i don't want to break the rules or something

OpenStudy (shiraz14):

You're welcome. Glad to be of assistance.